OpenStudy (anonymous):

Which of the following is a solution to the equation log 4 x + log 4 (x – 3) = 1 ?

OpenStudy (anonymous):

x = 6/5 x = 2 x = –3 x = 4

OpenStudy (anonymous):

log 4 x + log 4 (x – 3) = 1

OpenStudy (anonymous):

I never said the base was 10

OpenStudy (anonymous):

where did the 10 come from ?

OpenStudy (whpalmer4):

\[\log_4 x-\log_4(x-3) =1\]

OpenStudy (anonymous):

why is there minus signs?

OpenStudy (whpalmer4):

Oops that is + not - Adding logs is same as multiplying and then taking log of product, so we can rewrite as\[\log_4 x(x-3)=1\] which means \[x(x-3)=4^1\]

OpenStudy (whpalmer4):

Solve for the positive solution of \[x^2-3x=4\]

OpenStudy (whpalmer4):

(Not negative solution because that would cause use to take log of a negative number in our original equation and that gives us a complex number)

OpenStudy (whpalmer4):

@spiderman10 what did you get for answer?